Income Or Output Y Consumption Expenditure         C Investment Expenditure       I Government Expenditure          G...

Income

Or

Output

Y

Consumption

Expenditure

        C

Investment

Expenditure

      I

Government

Expenditure

         G

Net export

Expenditure

       NX

$4,000

3,925

100

100

25

4,100

4,000

100

100

25

4,200

4,075

100

100

25

4,300

4,150

100

100

25

4,400

4,225

100

100

25

4,500

4,300

100

100

25

4,600

4,375

100

100

25

4,700

4,450

100

100

25

4,800

4,525

100

100

25

4,900

4,600

100

100

25

5,000

4,675

100

100

25

b) Calculate expenditure multiplier
Answer #1

Marginal propensity to consume = Change in the consumer / Change in the income.

We will pick any two successive number for calculating that.

Old income 4400, new income 4500. Old consumption 4225, new consumption 4300.

Change in the income 4500 - 4400 = 100. Change in the consumption = 4300 - 4225 = 75.

75 / 100 = 0.75. So, the marginal propensity to consume in the economy is 0.75.  

After having the MPC we can calculate the expenditure multiplier = 1 / 1 - MPC   

= 1 / 1 - 0.75

= 1 / 0.25

= 4.

The expenditure multiplier in the economy is 4.